OpenStudy (anonymous):

air is being pumped into a spherical balloon at a rate of 2 cubic meters per minute. Find the rate of change of the radius when the radius is 3 meters. v=4/3pir^3

OpenStudy (anonymous):

were given the change of the volume V as a change in time. dV/dt = 2 m^3/min if we take the derivative of Volume d/dc (V) = d/dx (3piR^3) dV/dt = 4piR^2 dR/dt R = 3 m 2 = 4pi(3)^2 dR/dt 2 = 36pi dR/dt dR/dt = 1/18pi